# CHIPS-FF: Evaluating Universal Machine Learning Force Fields for Material Properties

### v3

@@ 3 Acknowledgments
All authors thank the National Institute of Standards and Technology for funding, computational, and data-management resources. This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose. The authors would like to acknowledge John Bonini for a fruitful discssion regarding the manuscript and software package.

### v4

@@ 2 Conflicts of Interest
3 Supporting Information
Description of software used, detailed description of test set, results for ExpCellFilter, CHIPS-FF tutorial
@@ 4 Acknowledgments
All authors thank the National Institute of Standards and Technology for funding, computational, and data-management resources. This work was performed with funding from the CHIPS Metrology Program, part of CHIPS for America, National Institute of Standards and Technology, U.S. Department of Commerce. Certain commercial equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identifications are not intended to imply recommendation or endorsement by NIST, nor it is intended to imply that the materials or equipment identified are necessarily the best available for the purpose. The authors would like to acknowledge John Bonini (NIST), Orbital Materials, and Preferred Networks for fruitful discussions regarding the manuscript and software package.